A portfolio consists of $15,400 in Stock M and $23,900 invested in Stock N. The expected return on these stocks is 9.20 percent and 12.60 percent, respectively. What is the expected return on the portfolio?
Answer #1

Total investment=(15400+23900)=$39300

Expected return of portfolio=Respective return*Respective investment weight

=(15400/39300*9.2)+(23900/39300*12.6)

which is equal to

=11.27%(Approx).